1. If you dine out at a restaurant and the bill come to R289,00 calculate the full bill if there was a discount of 15% included.
2. What fraction of an hour is 15minutes, 30 minutes, 45 minutes, 20 minutes and 10 minutes

Answer by Cromlix(4381)

Hi there,
100% - 15% = 85%
85% = R289.00
1% = R289.00/85
100% = R289.R34000/85 x 100
100% = R340
This is the full bill amount.
............
15/60 = 1/4 of an hour
30/60 = 1/2 of an hour
45/60 = 3/4 of an hour
20/60 = 1/3 of an hour
10/60 = 1/6 of an hour.
Hope this helps:-)